### 07 Mar 2024 v1

# New quasi-Einstein metrics on a two–sphere

Abstract.
We construct all axi-symmetric non–gradient $m$–quasi–Einstein structures on a two–sphere. This includes the spatial cross-section of the extreme Kerr black hole horizon corresponding to $m=2$, as well as a family of new regular metrics with $m\neq 2$ given in terms of hypergeometric functions. We also show that in the case $m=-1$ with vanishing cosmological constant the only orientable compact solution in dimension two is the flat torus, which proves that there are no compact surfaces with a metrisable affine connection with skew Ricci tensor.
1. Introduction
A Riemannian manifold $(M,g)$ of dimension $n$ is said to be $m-$quasi-Einstein if there exists a vector field $X\in\mathfrak{X}(M)$ and constant $\lambda$ such that
$\mbox{Ric}(g)=\frac{1}{m}X^{\flat}\otimes X^{\flat}-\frac{1}{2}{\mathcal{L}}_{X}g+\lambda g$ (1.1)
[…]
For generic $m$ a number of general results which constrain the topology and geometry of quasi-Einstein manifolds have been obtained in the case when the vector field $X$ is a gradient [3] (see also [4] if $m=2$). In particular all solutions to (1.1) on a two–sphere $S^{2}$ are trivial in this case. If $X$ is not a gradient and $m=2$, then the most general solution to (1.1) on $S^{2}$ belongs to the explicit two–parameter family of solutions arising from the extreme Kerr black hole horizon, which are parametrised by the mass and the cosmological constant  [10]. These solutions are all axi-symmetric.
In this paper we ask the simple question: are there any non-Einstein examples of $m-$quasi-Einstein manifolds on $M=S^{2}$ for $m\neq 2$? To the best of our knowledge no such example has been written down in the literature. Our main result is the following.
Theorem 1.1.
Let $(g,X)$ be a solution to the $m-$quasi–Einstein equation (1.1) on a two–dimensional connected surface $M$, with $\mathrm{d}X^{\flat}$ not identically zero, and a $U(1)$ isometric action.
  • Locally there exist coordinates $(x,\phi)$ on $M$, and a function $B=B(x)$ such that the quasi–Einstein structure is of the form
    $g=B(x)^{-1}\mathrm{d}x^{2}+B(x)\mathrm{d}\phi^{2},\quad X^{\flat}=\frac{-m}{x^{2}+1}\Big{(}x\mathrm{d}x-B(x)\mathrm{d}\phi\Big{)},$ (1.2)
    with
    $B(x)=\begin{cases}bx(x^{2}+1)^{-m/2}+c(x^{2}+1)^{-m/2}F(x)-\frac{\lambda(x^{2}+1)}{m+1}&\text{for $m\neq-1$}\\ x\Big{(}b-\lambda\operatorname{arcsinh}(x)\Big{)}\sqrt{x^{2}+1})+c(x^{2}+1)&\text{for $m=-1$}\end{cases}$ (1.3)
    where $b,c$ are constants and
    $F(x)\equiv_{2}F_{1}\Big{(}-\frac{1}{2},-\frac{m}{2},\frac{1}{2},-x^{2}\Big{)}$ (1.4)
    is the hypergeometric function.
  • If $b=0$ and
    $\lambda=0$ $\lambda>0$ $\lambda<0$
    $m>0$ $c>0$ $c>\frac{\lambda}{m+1}$ $c>\frac{|\lambda|}{m+1}c_{0}$
    $m\in(-1,0)$ $c>\frac{\lambda}{m+1}$
    $m=-1$ $c>0$
    $m<-1$ $c\in\Big{(}\frac{\lambda}{m+1},0\Big{)}$
    where
    $c_{0}=\min_{x>x_{0}}\frac{(x^{2}+1)^{\frac{m}{2}+1}}{|F(x)|}$
    and $x_{0}$ is the unique positive zero of $F$, then (1.2) smoothly extends to $S^{2}$.
  • Conversely all solutions to (1.1) on $S^{2}$ with a $U(1)$ isometric action arise from (1.2) with (1.3) and $b=0$, together with the restrictions on $c$ given in the above table.
Therefore $b=0$ together with the restriction on $c$ given in the second part of the Theorem are necessary and sufficient conditions for the structure (1.2) to extend to $S^{2}$. In the special case $m=2$ a much stronger result has been established [10]: the existence of a $U(1)$ isometry preserving $X$ follows from the global assumption that $M=S^{2}$. In this case (1.4) reduces to $F(x)=1-x^{2}$.
To arrive at the local normal form (1.2), (1.3) we shall use a reformulation (Proposition 2.4 in §2.4) of (1.1) as a fourth order non–linear PDE for the Kähler potential for the metric $g$. If $g$ admits a Killing vector, then either $[K,X]=0$, or $g$ is a metric of constant curvature (Proposition 3.1 in §3). In this axi-symmetric case the PDE for the Kähler potential reduces to an ODE solvable in terms of the hypergeometric functions (Theorem 3.3 in §3). To ensure that the solutions (1.2), (1.3) extend to smooth metrics on $S^{2}$, the function $B$ in (1.3) and its derivatives must satisfy certain regularity conditions at the end points (Lemma 4.1 in §4). Imposing these conditions under the additional assumption that $B$ is an even function leads to the proof of the second part of Theorem 1.1. The last part of the Theorem asserts that the even parity of $B$ is necessary as well as sufficient (Lemma 4.5 in §4). In particular we deduce that for all combinations of $(m,\lambda)$ allowed by the Gauss–Bonnet theorem (Proposition 2.1) there exists an open set $U\subset\mathbb{R}$ such that the quasi–Einstein structure corresponding to (1.3) extends to $S^{2}$ iff $c\in U$.
While the paper focuses on the genus zero case, some non–existence results can be established for other topologies. This has been done in [8] if $m=2$, and more recently [5] for general $m$. In §5 we shall use an old result of Milnor to establish the following.
Theorem 1.2.
Let $(g,X)$ satisfy the quasi–Einstein equation (1.1) with $\lambda=0$ and $m=-1$ on a compact orientable surface $M$. Then $X=0$ and $(M,g)$ is the flat torus.
In the Appendix we shall consider the general dimension $n$ and show that, if $m=2$ and $M$ is compact, then the Killing vector of $g$ proven to exist in [10] must preserve $X$. This strengthens the results of [10], where this inheritance property has been established only for $n=2$, or for general $n$ but $\lambda\leq 0$. In the rest of the paper with the exception of Lemma 2.3, some considerations in §5, and Theorem A.0 in the Appendix, we shall assume that $n=\dim{M}=2$.
Acknowledgements
MD is grateful to Jan Deredziński for helpful discussions. JL acknowledges support by the Leverhulme Research Project Grant RPG-2019-355. HK acknowledges the support of the NSERC Grant RGPIN-2018-04887. AC acknowledges the support of the Cambridge International Scholarship.
2. General properties of the quasi–Einstein equations on surfaces
In this section we shall establish some properties of (1.1) needed in the rest of the paper.
2.1. Topology
If $M$ is compact and without boundary, then its topology is constrained by the sign of $m$ and $\lambda$.
Proposition 2.1.
Let ${\tt g}_{M}$ be the genus of a two–dimensional closed orientable surface $M$, and let $(M,g,X)$ satisfy (1.1).
  • If $m>0$ and $\lambda>0$, then ${\tt g}_{M}=0$.
  • If $m>0$ and $\lambda=0$, then ${\tt g}_{M}\leq 1$ with an equality iff $(M,g)$ is the flat torus.
  • If $m<0$ and $\lambda<0$, then ${\tt g}_{M}>1$.
  • If $m<0$ and $\lambda=0$, then ${\tt g}_{M}\geq 1$ with an equality iff $(M,g)$ is the flat torus.
Proof.
Taking the trace of (1.1), and integrating the resulting equation over $M$, dropping the divergence term, and using the Gauss-Bonnet formula gives
$\int_{M}\Big{(}\frac{|X|^{2}}{m}+2\lambda\Big{)}\mbox{vol}_{M}=8\pi(1-{\tt g}_{M})\;,$ (2.1)
where $\mbox{vol}_{M}$ is the volume form of $(M,g)$. The statement of the proposition now follows immediately. In the genus one case with $\lambda=0$ we get $|X|^{2}=0$, which in the Riemannian signature implies that $X=0$. Substituting this back into (1.1) gives that $g$ is flat. ∎
2.2. Prolongation
The quasi–Einstein equations (1.1) do not form a closed system, as only the symmetrised part of the covariant derivative of $X^{\flat}$ is specified. It is however a system of finite type which means that adding additional unknowns, and cross–differentiating eventually determines all derivatives (see e.g. Appendix C in [9]). This process known as the prolongation has been carried over for (1.1) in [19], and it turns out that only one differentiation is needed.
[…]
$\nabla X^{\flat}=\frac{1}{m}X^{\flat}\otimes X^{\flat}+(\lambda-\tfrac{1}{2}R)g+\frac{1}{2}\Omega\mbox{vol}_{M}.$ (2.2)
Applying $\nabla$ to (2.2) and commuting the derivatives of $X$ leads to
$\text{d}\Omega=\frac{3}{m}\Omega X^{\flat}+{*}\text{d}R+\frac{1}{m}\left(2\lambda-\left(m+1\right)R\right){*}X^{\flat}\;,$ (2.3)
where $R$ is the scalar curvature and $\star$ the Hodge star operator of $g$. The system is now closed after one prolongation. The first constraint comes from setting the skew part of $\nabla^{2}\Omega$ to zero. It gives
$0=-\Delta R+\left(1+\frac{4}{m}\right)\langle X,\text{d}R\rangle+\frac{3}{m}\Omega^{2}+\frac{1}{m^{2}}(2\lambda-(m+1)R)(2|X|^{2}-2\lambda m+mR)\;,$ (2.4)
where $\Delta$ is the Laplace-Beltrami operator. Differentiating this and removing all derivatives of $X$ and $\Omega$ using (2.2) and (2.3) yields two more constraints
$0=$ $-\text{d}(\Delta R)+\frac{1}{m}\left(1+\frac{4}{m}\right)\langle X,\text{d}R\rangle X^{\flat}+\left(1+\frac{4}{m}\right)\text{Hess}(R)(X)+\left(\frac{1}{2}+\frac{8}{m}\right)\Omega\>{*}\text{d}R$
$\left[-\frac{2}{m}\left(1+\frac{1}{m}\right)|X|^{2}-\left(\frac{5}{2}+\frac{4}{m}\right)R+\lambda\left(3+\frac{8}{m}\right)\right]\text{d}R+\frac{18}{m^{2}}\Omega^{2}X^{\flat}$
$+\frac{4}{m^{2}}\left(2\lambda-(m+1)R\right)\left(\left(\lambda-\tfrac{1}{2}R\right)X^{\flat}+2\Omega\>{*}X^{\flat}+\frac{1}{m}|X|^{2}X^{\flat}\right).$ (2.5)
Remark 2.2.
In the constant curvature case (with $X$ non-identically zero) equations (2.3) and (2.5) imply $\Omega=0$ and $2\lambda=(m+1)R$ if $m\neq-1$ (if $m=-1$, then $R$ can be arbitrary but $\lambda$ must vanish).
2.3. Analyticity
We will need to refer to the real analyticity of (1.1) in our proof of Proposition 3.1 when $n=2$. The proof of the following result is however valid for any $n$. It follows the considerations by Pedersen and Tod for the Einstein-Weyl equation [20].
Lemma 2.3.
Let $(M,g,X)$ be a quasi-Einstein manifold of dimension $n$. In harmonic coordinates the components of $X$ and $g$ are real analytic.
Proof.
The contracted Bianchi identity applied to the quasi-Einstein equations (1.1) implies (see Lemma 2.1 in [1])
$\Delta X^{\flat}=-\lambda X^{\flat}+\left(\frac{2}{m}+\frac{1}{2}\right)\nabla_{X}X^{\flat}-\left(\frac{1}{m}-\frac{1}{4}\right)\nabla(|X|^{2})-\frac{1}{m}|X|^{2}X^{\flat}+\frac{2}{m}(\text{div }X)X^{\flat}.$ (2.6)
The only terms in (1.1) and (2.6) that contain second derivatives are Ric$(g)$ and $\Delta X^{\flat}$. In a harmonic coordinate chart $(U,(x^{i}))$ on $M$ we can write
$\text{Ric}(g)_{ij}=-\frac{1}{2}g^{kl}\frac{\partial^{2}g_{ij}}{\partial x^{k}\partial x^{l}}+\dots\hskip 11.38092pt\text{and}\hskip 11.38092pt\Delta X_{i}^{\flat}=g^{kl}\frac{\partial^{2}X_{i}^{\flat}}{\partial x^{k}\partial x^{l}}+\dots,$ (2.7)
where the dots denote terms involving at most first derivatives of components of $g$ and $X^{\flat}$. Note that the coordinate expression for the Laplace-Beltrami operator acting on one-forms does contain second derivatives of the metric, but these can be eliminated using the harmonic coordinate condition $g^{ij}\Gamma^{k}_{ij}=0$ and the expression for $g^{kl}\partial_{k}\partial_{l}g_{ij}$ coming from (1.1). Since the metric is positive definite, it follows that in harmonic coordinates (1.1) and (2.6) constitute an elliptic system of $\frac{1}{2}n(n+3)$ differential equations in $\frac{1}{2}n(n+3)$ unknowns $g_{ij}$ and $X_{k}^{\flat}$. Following the arguments of Deturck and Kazdan [7] we deduce that the components of $g$ and $X$ are real analytic in these coordinates. ∎
2.4. Kähler potential
In this section we shall reformulate (1.1) as a scalar PDE for a Kähler potential. In [10] this formulation was used in the particular case when $m=2$. We refer the reader to [10] for the general discussion of Kähler potentials and their global properties on $S^{2}$.
Proposition 2.4.
Locally there exist complex coordinates $(\zeta,\bar{\zeta})$ and a Kähler potential $f$ on $M$ such that
[…]
and the system (1.1) reduces to a single 4th order PDE for the Kähler potential $f$
$\frac{2}{m}(f_{\zeta}f_{\bar{\zeta}})^{2}(f_{\zeta\zeta\bar{\zeta}\bar{\zeta}}f_{\zeta\bar{\zeta}}-f_{\zeta\zeta\bar{\zeta}}f_{\zeta\bar{\zeta}\bar{\zeta}})+\frac{4\lambda}{m}(f_{\zeta\bar{\zeta}})^{3}(f_{\zeta})^{2}(f_{\bar{\zeta}})^{2}$ (2.9)
$-(f_{\zeta\bar{\zeta}})^{3}(f_{\bar{\zeta}\bar{\zeta}}(f_{\zeta})^{2}+f_{\zeta\zeta}(f_{\bar{\zeta}})^{2})+(f_{\zeta\bar{\zeta}})^{2}(f_{\zeta}(f_{\bar{\zeta}})^{2}f_{\zeta\zeta\bar{\zeta}}+f_{\bar{\zeta}}(f_{\zeta})^{2}f_{\zeta\bar{\zeta}\bar{\zeta}})+2(f_{\zeta\bar{\zeta}})^{4}f_{\zeta}f_{\bar{\zeta}}=0.$
Proof.
Choose holomorphic coordinates and an associated Kähler potential so $g$ is of the form (2.8). In these coordinates write
$X^{\flat}=A\text{d}\zeta+\overline{A}\text{d}\bar{\zeta},$
where $A$ is a smooth complex valued function. The $(\zeta\zeta)$ and $(\bar{\zeta}\bar{\zeta})$ components of (1.1) yield
$\partial_{\zeta}A-\frac{1}{m}A^{2}-\frac{f_{\zeta\zeta\bar{\zeta}}}{f_{\zeta\bar{\zeta}}}A=0,$
and the complex conjugate of this equation. These two equations can be integrated to find $A$ in terms of $f$ and its derivatives to be
$A=-m\frac{f_{\zeta\bar{\zeta}}}{\partial_{\bar{\zeta}}f+\partial_{\bar{\zeta}}\overline{C}},$ (2.10)
where $C=C(\zeta)$ is an arbitrary holomorphic function. The freedom of adding a holomorphic function and its conjugate to $f$ can be used to set $C$ to zero. This choice yields the form of $X^{\flat}$ in (2.8). Finally, using the formula for the Ricci scalar
$R=-(\Delta_{0}f)^{-1}\Delta_{0}(\ln{(\Delta_{0}f)}),\quad\mbox{where}\quad\Delta_{0}=4\frac{\partial^{2}}{\partial\zeta\partial\bar{\zeta}}$ (2.11)
is the flat Laplacian, the $(\zeta\bar{\zeta})$ component of (1.1) gives the 4th order PDE (2.9) for $f$. ∎
If the gradient and Laplacian are taken with respect to the Cartesian coordinates $(\frac{1}{2}(\zeta+\bar{\zeta}),\frac{1}{2i}(\zeta-\bar{\zeta}))$ then, expanding the derivatives and isolating the numerator in (2.9), we find that this equation is equivalent to a vector PDE on $\mathbb{R}^{2}$ given by
$\frac{1}{m}{\Delta_{0}}^{2}f-\frac{1}{m}\frac{|\nabla(\Delta_{0}f)|^{2}}{\Delta_{0}f}+2\frac{(\Delta_{0}f)^{3}}{|\nabla f|^{2}}+\Delta_{0}f\nabla f\cdot\nabla\Big{(}\frac{\Delta_{0}f}{|\nabla f|^{2}}\Big{)}+\frac{2}{m}\lambda(\Delta_{0}f)^{2}=0.$ (2.12)
Here $\nabla$ and $\Delta_{0}=\nabla\cdot\nabla$ are the nabla operator and the Laplacian of the flat Euclidean metric.
3. Symmetry reduction of the quasi–Einstein equation
In this section we shall show that if $g$ is assumed to be axisymmetric then the 4th order ODE resulting from the PDE (2.9) is solvable by quadratures.
3.1. Inheritance of symmetry
We shall first demonstrate that if metric admits a Killing vector field, the constraints (2.4) and (2.5) resulting from the prolongation procedure can be used to deduce that $X$ inherits the symmetry from $g$.
Proposition 3.1.
Let $(g,X)$ be a solution to the quasi-Einstein equations (1.1) on a two-dimensional connected surface $M$ admitting a Killing vector field $K$. Then either $[K,X]=0$ or $g$ has constant curvature.
Proof.
[…]
We start by solving (2.4) for $X_{\theta}^{2}$ to find
$X_{\theta}^{2}=k^{2}\left(\frac{m^{2}\Delta R-m(m+4)X_{r}R^{\prime}-3m\Omega^{2}}{2(2\lambda-(m+1)R)}+m\lambda-\frac{1}{2}mR-X_{r}^{2}\right),$ (3.1)
where the prime denotes a derivative with respect to $r$. The $r$- and $\theta$-components of (2.5) are
00 0 $=-(\Delta R)^{\prime}+\tfrac{m+4}{m^{2}}X_{r}^{2}R^{\prime}+\tfrac{m+4}{m}X_{r}R^{\prime\prime}+R^{\prime}\left(-\tfrac{2(m+1)}{m^{2}}(X_{\theta}^{2}k^{-2}+X_{r}^{2})-(\tfrac{5}{2}+\tfrac{4}{m})R+\lambda(3+\tfrac{8}{m})\right)$
$\hskip 11.38092pt+\tfrac{18}{m^{2}}\Omega^{2}X_{r}+\tfrac{4}{m^{2}}(2\lambda-(m+1)R)\left(X_{r}(\lambda-\tfrac{1}{2}R)+2\Omega k^{-1}X_{\theta}+\tfrac{1}{m}X_{r}(X_{r}^{2}+k^{-2}X_{\theta}^{2})\right),$ (3.2)
00 0 $=R^{\prime}\left(\tfrac{m+4}{m^{2}}X_{\theta}X_{r}+\tfrac{m+4}{m}X_{\theta}k^{\prime}k^{-1}-\tfrac{m+16}{2m}\Omega k\right)+\tfrac{18}{m^{2}}\Omega^{2}X_{\theta}$
$\hskip 11.38092pt+\tfrac{4}{m^{2}}(2\lambda-(m+1)R)\left(X_{\theta}(\lambda-\tfrac{1}{2}R)-2\Omega kX_{r}+\tfrac{1}{m}X_{\theta}(X_{r}^{2}+k^{-2}X_{\theta}^{2})\right).$ (3.3)
We use (3.1) to eliminate the $X_{\theta}^{2}$ terms in (3.2) and (3.3). After doing this there is only one occurrence of $X_{\theta}$ left in (3.2) in a term proportional to $X_{\theta}\Omega$. We can solve this for $X_{\theta}\Omega$ and use it to eliminate $X_{\theta}$ completely from (3.3) times $\Omega$. The resulting equation is polynomial in $u=\Omega^{2}$ and $Y=(m+4)R^{\prime}X_{r}$. It can be written in the form
$P_{1}(u,Y)=\alpha_{0}(Y)+\alpha_{1}(Y)u+\alpha_{2}(Y)u^{2}=0,$ (3.4)
where $\alpha_{0},\alpha_{1},\alpha_{2}$ is a polynomial of degree $3,2,1$ with leading term $\frac{1}{8}Y^{3},-3Y^{2},18Y$ respectively. The lower order coefficients are complicated expressions involving up to 2 derivatives of $k$ and up to 3 derivatives of $R$.
We obtain a second polynomial constraint in $u$ and $Y$ by contracting (2.5) with $X$ and using (3.1) and the expression for $X_{\theta}\Omega$ obtained from (3.2) to eliminate any occurrences of $X_{\theta}$. This leads to
$P_{2}(u,Y)=\beta_{0}(Y)+\beta_{1}(Y)u-6u^{2}=0,$ (3.5)
where $\beta_{0},\beta_{1}$ is a polynomial of degree $2,1$ with leading term $\beta_{1}=-\frac{1}{4}Y+\dots$ and
$\beta_{0}=\left[\frac{1}{6}+\frac{5m}{16(m+4)}+\frac{2\lambda-(m+1)R}{3(m+4)R^{\prime}}\left(\frac{R^{\prime\prime}}{R^{\prime}}-\frac{k^{\prime}}{k}\right)\right]Y^{2}+\dots.$ (3.6)
The lower order terms again contain $k,k^{\prime},k^{\prime\prime}$ and up to 3 derivatives of $R$. Let us view $P_{1}$ and $P_{2}$ as polynomials in $u$. The resultant res$(P_{1},P_{2})$ is a polynomial in $Y$ of degree at most 6 with coefficients depending on $r$ only. If there is a solution such that $X_{r}$ (and hence $Y$) depends on $\theta$, this polynomial must be identically zero. To calculate the coefficient $\gamma_{6}$ of $Y^{6}$ we only need the leading coefficients of $\alpha_{i}$ and $\beta_{i}$. Setting $\gamma_{6}$ to zero gives
$R^{\prime\prime}=\frac{k^{\prime}R^{\prime}}{k}-\frac{5(m+1)(R^{\prime})^{2}}{4(2\lambda-(m+1)R)}.$ (3.7)
For the coefficient $\gamma_{5}$ of $Y^{5}$ we need the first subleading terms in $\alpha_{i}$ and $\beta_{i}$. Requiring $\gamma_{5}$ to vanish imposes
$\left(R^{\prime\prime}-\frac{k^{\prime}R^{\prime}}{k}+\frac{5(m+1)(R^{\prime})^{2}}{4(2\lambda-(m+1)R)}\right)\delta-\frac{25}{3}(m+4)^{2}(R^{\prime})^{3}(2\lambda-(m+1)R)=0,$ (3.8)
where
$\delta=\>$ $m(23m+68)(m+4)(R^{\prime})^{2}k^{\prime}k^{-1}-15m(m+4)^{2}R^{\prime}R^{\prime\prime}+32m(2\lambda-(m+1)R)(\Delta R)^{\prime}$
$-16(3m+4)(m+1)(m+4)(R^{\prime})R^{2}+32(5m^{2}+34m+32)\lambda R(R^{\prime})+48m(m+1)R^{\prime}\Delta R$
$-128(m+8)\lambda^{2}R^{\prime}-\tfrac{75}{4}m(m+1)(m+4)^{2}(2\lambda-(m+1)R)^{-1}(R^{\prime})^{3}.$ (3.9)
Thus given our assumptions the coefficients of $Y^{5}$ and $Y^{6}$ can not vanish simultaneously. Hence res$(P_{1},P_{2})$ is not identically zero, and so we must have $\partial_{\theta}X_{r}=0$. But now (3.5) is an equation for $u$ with coefficients only depending on $r$, implying $\partial_{\theta}u=0$. From (3.1) we then also find $\partial_{\theta}X_{\theta}=0$, and therefore $[K,X]=0$.
It remains to consider the cases $(m,\lambda)=(-1,0)$ and $m=-4$. If $(m,\lambda)=(-1,0)$ we can use (2.4) to eliminate the $\Omega^{2}$ term in (3.2). The resulting equation is
$0=-(\Delta R)^{\prime}+\frac{3}{2}RR^{\prime}-3X_{r}(R^{\prime\prime}+2\Delta R)-15X_{r}^{2}R^{\prime}.$ (3.10)
It follows that $\partial_{\theta}X_{r}=0$ because the coefficients of (3.10) depend on $r$ only and are not all zero. With the same argument as before equation (2.4) implies $\partial_{\theta}\Omega=0$ and (3.3) gives $\partial_{\theta}X_{\theta}=0$ unless $\Omega=0$, in which case we see from (2.3) that the curvature is constant.
If $m=-4$ we can use (2.4) to express $|X|^{2}$ in terms of $\Omega^{2}$. Using this the equations (2.5) become linear in $X^{\flat}$. They can be solved for $X^{\flat}$ in terms of $\Omega$, giving an expression of the form
$\rho(\Omega)X^{\flat}=\sigma(\Omega),$ (3.11)
where $\rho$ is a non-negative function and $\sigma$ is a 1-form (these are also functions of $k,R$ are their derivatives). The function $\rho$ and (the components of) $\sigma$ are polynomials of degree up to $4$ in $\Omega$ with leading terms $\Omega^{4}$ and $-3(2\lambda+3R)^{-1}(\text{d}R)\Omega^{4}$ respectively. The coefficients of $\rho,\sigma_{r}$ and $\sigma_{\theta}$ depend on $r$ only. Taking the norm of (3.11) and plugging this into (2.4) multiplied by $\rho(\Omega)^{2}$ gives a polynomial constraint in $\Omega$ with leading term $-\frac{3}{4}\Omega^{10}$ coming from the $\frac{3}{m}\Omega^{2}$ term in (2.4). It follows that $\partial_{\theta}\Omega=0$. Finally, if $\Omega\neq 0$ then $\rho>0$ and hence Lie deriving (3.11) along $K$ shows that $\mathcal{L}_{K}X^{\flat}=0$, whereas if $\Omega=0$ this follows by Lie deriving (2.3) along $K$.
We conclude that whenever $R$ is not constant we have $[K,X]=0$.∎
Remark 3.2.
In the constant curvature case $X^{\flat}$ is closed (see Remark 2.2). Then, locally there is always a Killing vector $K$ commuting with $X$. If locally $X^{\flat}=\text{d}h$, then $K^{\flat}={*}\text{ d}(e^{-h/m})$.
3.2. General axi-symmetric solution
We are now ready to present the local normal form of the most–general axi-symmetric quasi-Einstein structure in two dimensions.
Theorem 3.3.
If a solution $(M,g,X)$ to the quasi–Einstein equation (1.1) on a two-dimensional surface $M$ admits an axial Killing vector field $K$, then there exist coordinates $(x,\phi)$ on $M$, and a function $B=B(x)$ such that the quasi–Einstein structure is of the form
$g=B(x)^{-1}\mathrm{d}x^{2}+B(x)\mathrm{d}\phi^{2},\quad X^{\flat}=\frac{-m}{x^{2}+\beta^{2}}\Big{(}x\mathrm{d}x-B(x)\beta\mathrm{d}\phi\Big{)},$ (3.12)
[…]
$B(x)=\begin{cases}bx(x^{2}+\beta^{2})^{-m/2}+c\;F(x/\beta)(x^{2}+\beta^{2})^{-m/2}-\frac{\lambda(x^{2}+\beta^{2})}{m+1}&\text{for $m\neq-1,\beta\neq 0$}\\ x\Big{(}b-\lambda\operatorname{arcsinh}(x/|\beta|)\Big{)}\sqrt{x^{2}+\beta^{2}})+c(x^{2}+\beta^{2})&\text{for $m=-1,\beta\neq 0$}\\ bx^{1-m}+c-\frac{\lambda x^{2}}{m+1}&\text{for $m\neq\{-1,1\},\beta=0$}\\ b\ln{x}+c-\frac{\lambda x^{2}}{2}&\text{for $m=1,\beta=0$}\\ cx^{2}+b-\lambda x^{2}\ln{x}&\text{for $m=-1,\beta=0$.}\end{cases}$ (3.13)
Here $b,c$ are arbitrary constants and $F(x):=_{2}F_{1}(-\frac{1}{2},-\frac{m}{2},\frac{1}{2},-x^{2})$ is the hypergeometric function.
Proof.
Recall Proposition 3.1 guarantees that $[K,X]=0$. We shall first establish that the coordinates can be chosen so that $K(f)=\mbox{const}$, where $f$ is the Kähler potential from Proposition 2.4. We can rewrite $X^{\flat}$ in (2.8) in the form
$X^{\flat}=-\tfrac{1}{m}|X|^{2}\text{d}f.$ (3.14)
Therefore, $\mathcal{L}_{K}X=0$ implies invariance of $\text{d}f$ along a Killing field $K$. Conversely, we can invert the previous relation so $X^{\flat}=-m\;\text{d}f/|\text{d}f|^{2}$, which implies that if $\text{d}f$ is Lie-derived by a Killing field $K$ then so is $X^{\flat}$. We conclude that $[K,X]=0$ iff ${\mathcal{L}}_{K}(\text{d}f)=\text{d}(K(f))=0$ or equivalently if $K(f)$ is a constant.
Choose the holomorphic coordinate $\zeta=re^{i\phi}$ on $M$ (so that the $U(1)$-action is $\zeta\rightarrow e^{i\psi}\zeta$), we see that the metric takes the form (2.8) and the Kähler potential satisfies
$K(f_{\zeta\bar{\zeta}})=0,\quad\mbox{where}\quad K=i(\zeta\partial_{\zeta}-\bar{\zeta}\partial_{\bar{\zeta}}).$ (3.15)
The solution to this is
$f=\rho(r)+R(\zeta)+\bar{R}(\bar{\zeta}),\quad\mbox{where}\quad r=|\zeta|,$
and $\rho(r)$, $R(\zeta)$, $\bar{R}(\bar{\zeta})$ are integration functions. Moreover $[K,X]=0$ iff $K(f)=\mbox{const}$ which gives
$\zeta R_{\zeta}=\frac{1}{2}(\alpha+i\beta)$
for some constants $(\alpha,\beta)$.
Now set $r=e^{s}$ and $x=d\rho/ds$, to find that
$g=\frac{\text{d}x}{\text{d}s}\Big{(}\text{d}s^{2}+\text{d}\phi^{2}),\quad X^{\flat}=-\frac{m}{4}\frac{\text{d}x}{\text{d}s}\Big{(}\frac{\text{d}s+i\text{d}\phi}{x/2+\bar{\zeta}\bar{R}_{\bar{\zeta}}}+\frac{\text{d}s-i\text{d}\phi}{x/2+\zeta R_{\zeta}}\Big{)}.$
Setting
$B=\Big{(}\frac{\text{d}s}{\text{d}x}\Big{)}^{-1}$ (3.16)
puts $(g,X)$ in the form (3.12), where the constant $\alpha$ has been set to 00 by translating the function $x$. The PDE (2.9) reduces to an ODE which linearises under the reciprocal transformation (3.16). The resulting equation for $B=B(x)$ is
$B^{\prime\prime}+\frac{m}{(\beta^{2}+x^{2})^{2}}\Big{(}x(\beta^{2}+x^{2})B^{\prime}+2\beta^{2}B\Big{)}+2\lambda=0,\quad\mbox{where}\quad\beta=\mbox{const}$ (3.17)
which can be solved explicitly by (3.13). ∎
If $X^{\flat}$ is not closed, which requires $\beta\neq 0$, then without loss of generality we can set $\beta=1$. To see this, set $x=\beta{\tilde{x}},\phi=\beta^{-1}\tilde{\phi},B=\beta^{2}\widetilde{B}$. Rewriting (3.12) and (3.17) in terms of $(\widetilde{B},\tilde{x},\tilde{\phi})$ and dropping tilde has an effect of setting $\beta=1$, and reducing (3.13) to the form (1.3). However, with such a choice the coordinate $\phi$ is no longer $2\pi$-periodic.
Remark 3.4.
One can also derive the local form of the solution (3.12) and (3.13) by generalising the method used for the $m=2$ case in the context of vacuum near-horizon geometries [14, Theorem 4.3]. If $X^{\flat}$ is not closed, so we can set $\beta=1$, we can write the Killing field of (3.12) as $K^{\flat}=\Gamma X^{\flat}+\frac{m}{2}\text{d}\Gamma$ where $\Gamma=(1+x^{2})/m$ is a smooth positive function. This generalises the form of the Killing field for the $m=2$ case, that was proven must exist on $S^{2}$ [10].
4. Extension to a two-sphere
We will make use of the following standard result. We include a proof for completeness.
Lemma 4.1.
The metric (1.2) extends to a smooth metric on $S^{2}$ if and only if there exist adjacent simple zeros $x_{1}<x_{2}$ of $B(x)$ such that $B(x)>0$ for all $x_{1}<x<x_{2}$ and
$B^{\prime}(x_{1})=-B^{\prime}(x_{2})$ (4.1)
where $\phi\sim\phi+p$ is periodically identified with period $p=4\pi/|B^{\prime}(x_{i})|$.
Proof.
First note that any Killing field on $S^{2}$ must have closed orbits. The metric (1.2) is valid in a chart $(x,\phi)$ such that $B(x)>0$. In this chart $K=\partial_{\phi}$ is a Killing field and $|K|^{2}=B$. Since the orbits of $K$ are closed the coordinate must be periodically identified $\phi\sim\phi+p$ for some period $p>0$. Let $(x_{1},x_{2})$ be the maximal interval such that $B(x)>0$ for all $x\in(x_{1},x_{2})$ (including possibly an infinite or semi-infinite interval). Requiring that $g$ extends to a smooth metric on a compact surface $M$ implies that $(x_{1},x_{2})$ must be a finite interval, so by maximality and continuity of $|K|^{2}$ we must have $B(x_{1})=B(x_{2})=0$. Thus $|K|^{2}=B>0$ for $x\in(x_{1},x_{2})$ and $|K|^{2}\to 0$ as $x\to x_{1}$ or $x\to x_{2}$, so we may add points corresponding to $x=x_{1}$ and $x=x_{2}$ to obtain a compact surface $M$ (these could correspond to the same point). These points are zeros of the Killing field $K$.
For the fixed point $x=x_{1}$ we can introduce geodesic coordinates $(s,\phi)$ by $s:=\int_{x_{1}}^{x}\text{d}x/\sqrt{B(x)}$ in which the metric takes the form $g=\text{d}s^{2}+f(s)\text{d}\phi^{2}$ where $f(s):=B(x(s))$. In general, the metric has a conical singularity at $s=0$ which is absent iff, for closed constant-$s$ curves, the ratio of the circumference to radius approaches $2\pi$ as $s\to 0$, that is,
$2\pi=\lim_{s\to 0}\frac{p\sqrt{f(s)}}{s}=p\lim_{s\to 0}\frac{\text{d}}{\text{d}s}\sqrt{f(s)}=\frac{p}{2}B^{\prime}(x_{1})\qquad\implies\qquad p=\frac{4\pi}{B^{\prime}(x_{1})}\;,$ (4.2)
where $p$ is the period of an orbit of $K=\partial_{\phi}$. An identical argument can be applied to the fixed point $x=x_{2}$ by defining $s:=\int_{x}^{x_{2}}\text{d}x/\sqrt{B(x)}$ which shows that the conical singularity at $s=0$ is absent iff $p=-4\pi/B^{\prime}(x_{2})$.
The vector field $K$ has one or two simple zeros so $M=\mathbb{RP}^{2}$ or $S^{2}$ by the Poincaré-Hopf theorem. If we assume that $M$ is orientable, then it is diffeomorphic to $S^{2}$, which can independently be verified by the Gauss–Bonnet theorem: The scalar curvature is $R_{g}=-2B^{\prime\prime}(x)$ and hence
$\frac{1}{4\pi}\int_{M}R_{g}\text{vol}_{g}=-\frac{p}{4\pi}[B^{\prime}(x)]_{x_{1}}^{x_{2}}=2\;,$ (4.3)
so the surface is $M=S^{2}$. ∎
Remark 4.2.
In the case of $\mathbb{RP}^{2}$ the two zeros of $B$ are identified by the antipodal map, and correspond to the same point on $M$. In (4.3) the LHS has to be halved to take the $\mathbb{Z}_{2}$ quotient into account. In fact, all global solutions of Theorem 1.1 descend to $\mathbb{RP}^{2}$.
4.1. Hypergeometric function
To investigate the regularity conditions imposed by Lemma 4.1 on the local form (1.2) we will need a couple of results about the hypergeometric function.
Lemma 4.3.
The function $F(x):=_{2}F_{1}(-1/2,-m/2,1/2,-x^{2})$ admits an integral representation
$F(x)=1-x\int_{0}^{x}\Big{(}\frac{(y^{2}+1)^{\frac{m}{2}}-1}{y^{2}}\;\mathrm{d}y\Big{)}.$ (4.4)
It is monotonically increasing if $x>0$ and $m<0$ and monotonically decreasing for $x>0$ if $m>0$.
Proof.
The key observation is that for all $m\neq 0$
$\frac{\text{d}}{\text{d}x}\Big{(}x^{-1}F(x)\Big{)}=-x^{-2}(1+x^{2})^{m/2}.$ (4.5)
Moreover, from the hypergeometric series we have $F(x)=1+O(x^{2})$. Therefore integrating (4.5) gives the integral representation (4.4). Noting that $(y^{2}+1)^{m/2}$ is greater (smaller) than $1$ if $m$ is positive (negative) establishes the monotonicity claim. ∎
The representation (4.4) implies that the asymptotic form of $F(x)$ as $x\rightarrow\infty$ is
$F(x)\sim\begin{cases}\sqrt{\pi}\frac{\Gamma\Big{(}-\frac{m}{2}+\frac{1}{2}\Big{)}}{\Gamma\Big{(}-\frac{m}{2}\Big{)}}x&\text{for $m\in(-\infty,0)\cup(0,1)$}\\ -x\ln{x}&\text{for $m=1$}\\ -\frac{1}{m-1}x^{m}&\text{for $m>1$.}\end{cases}$ (4.6)
The coefficient of the linear term is negative for $m\in(0,1)$ and positive for $m<0$.
The next lemma shows the existence of two simple zeros of $F$ for $m>0$.
Lemma 4.4.
For $m>0$ the function $F$ is monotonically decreasing faster than a linear function: For any $x_{*}>0$ there exists a constant $a(x_{*})$ such that $F(x)<1-a(x_{*})x$ if $x>x_{*}$. In particular $F$ has precisely two simple zeros.
Proof.
It is convenient to define
$f(t)=\frac{(1+t^{2})^{m/2}-1}{t^{2}}$
which for $m>0$ is a strictly positive function for $t>0$. Then we can write $F(x)=1-x\int_{0}^{x}f(t)\text{d}t$. Note that $F(0)=1$ so $F(x)>0$ near $x=0$. Positivity of $f(t)$ implies that for $y>x>0$ we have $\int_{0}^{y}f(t)\text{d}t>\int_{0}^{x}f(t)\text{d}t$. Thus $F(x)$ for $m>0$ is monotonically decreasing for $x>0$. However, we need a little more. In particular, fix $x_{*}>0$, then for all $x>x_{*}$
$x\int_{0}^{x}f(t)\text{d}t>x\int_{0}^{x_{*}}f(t)\text{d}t$
and hence
$F(x)<1-x\int_{0}^{x_{*}}f(t)\text{d}t\;.$
Thus $F(x)$ for $m>0$ and $x>x_{*}$ is monotonically decreasing faster than a linear function.
Now assume for contradiction that $F(x)>0$ for all $x>0$. But then for all $x>x_{*}$ we have that $x\int_{0}^{x_{*}}f(t)\text{d}t<1$ which is a contradiction. Therefore it must be the case that there exists a zero of $F(x)$ at say $x=x_{2}>0$. Furthermore, monotonicity of $F(x)$ means that this is the only zero for $x>0$. Finally, recall that $F(x)$ is an even function and therefore it must have precisely two zeros at $x=\pm x_{2}$. The fact they are simple follows from evaluating (4.5) at the roots which gives
$F^{\prime}(x_{i})=-\frac{(1+x_{i}^{2})^{m/2}}{x_{i}}$
and so $F^{\prime}(x_{2})=-F^{\prime}(x_{1})<0$. ∎
4.2. The parity of $B$
The function $B$ appearing in the local normal form (1.2) is a sum of an even part given by the multiple of the hypergeometric function, and an odd part. We shall now show that the regularity conditions of Lemma 4.1 force $B$ to be even.
Lemma 4.5.
The metric (1.2) extends to a smooth metric on $M=S^{2}$ if and only if the function $B$ given by (1.3) is even (and so $b=0$).
Proof.
First observe that the function $B$ given by (1.3) satisfies the following first order ODE
$\frac{\text{d}}{\text{d}x}\left(\frac{B(x)(x^{2}+1)^{m/2}}{x}\right)=-\frac{(1+x^{2})^{m/2}}{x^{2}}(\alpha+\lambda x^{2}),\qquad\alpha:=\begin{cases}c-\frac{\lambda}{m+1}&\text{for $m\neq-1$}\\ c&\text{for $m=-1$}\end{cases}$ (4.7)
where to evaluate the derivative we used the identity (4.5). We also note that $B(0)=\alpha$.
Let $x_{1}<x_{2}$ be two adjacent zeros of $B$ such that $B^{\prime}(x_{1})=-B^{\prime}(x_{2})=\delta>0$. Assume first that $\lambda\neq 0$ and $m\neq-1$ and set
$\hat{F}(x):=\frac{B(x)(x^{2}+1)^{\frac{m}{2}}}{\lambda x}\;.$ (4.8)
Setting $\gamma=-\frac{B(0)}{\lambda}$ and using (4.7) shows that
$\hat{F}^{\prime}(x)=-\frac{(x^{2}+1)^{\frac{m}{2}}(x^{2}-\gamma)}{x^{2}}.$ (4.9)
If $x_{1}$ and $x_{2}$ are non-zero we can write
$\hat{F}^{\prime}(x_{1})=\frac{\delta(x_{1}^{2}+1)^{\frac{m}{2}}}{\lambda x_{1}}\hskip 5.69046pt\text{and}\hskip 5.69046pt\hat{F}^{\prime}(x_{2})=-\frac{\delta(x_{2}^{2}+1)^{\frac{m}{2}}}{\lambda x_{2}}.$ (4.10)
If $x_{1}$ equals zero the function $\hat{F}$ is well-defined on all of $\mathbb{R}$ and is strictly decreasing with $\hat{F}(0)=\frac{\delta}{\lambda}$ and $\hat{F}^{\prime}(x)=-(x^{2}+1)^{\frac{m}{2}}$. Equation (4.10) is still valid for $x_{2}$ and gives $x_{2}=\frac{\delta}{\lambda}$. Since $x_{2}$ has to be positive this already excludes $\lambda<0$. Moreover, depending on the sign of $m$, $\hat{F}^{\prime}(x)$ is either always greater than $-1$ or less than $-1$. In both cases we can not have $\hat{F}(x_{2})=0$. Similarly, if $x_{2}=0$ we have $\hat{F}(0)=-\frac{\delta}{\lambda}$ and $x_{1}=-\frac{\delta}{\lambda}$, but $\hat{F}$ decreases either too fast $(m>0)$ or too slow $(m<0)$ for this to happen. Hence $x_{1}$ and $x_{2}$ are both non-zero.
Eliminating $\delta$ between the two equations in (4.10) leads to
$\left(1-\frac{\gamma}{x_{1}x_{2}}\right)(x_{1}+x_{2})=0.$ (4.11)
There are now three cases to consider:
  • If $\lambda=0$ repeating the calculation above with $\hat{F}$ replaced by $\lambda\hat{F}$ shows that we must have $x_{1}=-x_{2}$, and $b=0$.
  • If $\lambda<0$ and $x_{1}=-x_{2}$ we find $b=0$ and the result follows, so let us instead assume $x_{1}x_{2}=\gamma$. If $\gamma<0$ we have $B(0)<0$ and the zeros have opposite sign, but $B$ is not positive between the zeros. If $\gamma>0$ the zeros have the same sign and from (4.10) we need $\hat{F}^{\prime}(x_{2})$ to have the same sign as $x_{2}$. Since $x_{1}<x_{2}$ we either have $x_{2}>\sqrt{\gamma}$ or $-\sqrt{\gamma}<x_{2}<0$. However, from (4.9) we find that $\hat{F}^{\prime}(x_{2})$ and $x_{2}$ have opposite signs which is a contradiction.
  • If $\lambda>0$, we again assume that $x_{1}x_{2}=\gamma$ as otherwise $x_{1}=-x_{2}$ and $b=0$. Define
    $G(x):=\hat{F}(x)-\hat{F}\left(\frac{\gamma}{x}\right).$ (4.12)
    Note that $G(x_{1})=G(x_{2})=0$. We have
    $G^{\prime}(x)=\hat{F}^{\prime}(x)+\frac{\gamma}{x^{2}}\hat{F}^{\prime}\left(\frac{\gamma}{x}\right)=\left(\frac{\gamma}{x^{2}}-1\right)\left((x^{2}+1)^{\frac{m}{2}}-\left(\frac{\gamma^{2}}{x^{2}}+1\right)^{\frac{m}{2}}\right).$ (4.13)
    The function $G^{\prime}(x)$ vanishes if and only if $x^{2}=|\gamma|$. If $\gamma>0$ we have $G(\pm\sqrt{\gamma})=0$. From (4.13) we know that $G$ is monotonic between $x_{2}$ and sgn$(x_{2})\sqrt{\gamma}$, contradicting the fact that $G(x_{2})=0$.
    If $\gamma<0$ we must have $x_{1}<0$ and $x_{2}>0$ and $B(0)>0$. We will show that $G$ has no zeros in this case.
    If $\gamma<0$ and $m>0$ the function $G$ has a maximum at $x=\sqrt{-\gamma}$. Since $x_{2}>0$ in order for $G(x_{2})=0$ to be possible we need $G(\sqrt{-\gamma})\geq 0$. To show that this is not the case, set $u:=\frac{c}{\lambda}>\frac{1}{m+1}$ and consider the function
    $H(u)$ $:=\frac{1}{2u}\left(\hat{F}(\sqrt{-\gamma})-\hat{F}(-\sqrt{-\gamma})\right)$ (4.14)
    $=\frac{1}{\sqrt{u-\tfrac{1}{m+1}}}\left(F\left(\sqrt{u-\tfrac{1}{m+1}}\right)-\frac{(u+1-\tfrac{1}{m+1})^{\frac{m}{2}+1}}{u(m+1)}\right).$
    Note that $H$ tends to zero as $u\to\frac{1}{m+1}$. A calculation using (4.5) shows
    $H^{\prime}(u)=-\frac{m(1+u-\tfrac{1}{m+1})^{\frac{m}{2}}\sqrt{u-\tfrac{1}{m+1}}}{(m+1)u^{2}}.$ (4.15)
    We see that $H$ is strictly decreasing, implying that $G(\sqrt{-\gamma})=2uH(u)<0$.
    If $\gamma<0$ and $m\in(-1,0)$ we find that $G$ has a minimum at $x=\sqrt{-\gamma}$, so in order for $G$ to have zeros we require $G(\sqrt{-\gamma})\leq 0$. However, this is not possible since $H$ is now increasing and therefore positive for $u>\frac{1}{m+1}$.
    If $\gamma<0$ and $m<-1$ the function $G$ again has a minimum at $\sqrt{-\gamma}$ and $H$ is decreasing. For $u<0$ this implies $H$ is negative and so $G(\sqrt{-\gamma})>0$. For $u=0$ direct evaluation shows $G(\sqrt{-\gamma})>0$. If $u$ is positive, so is $H$ because using (4.6) it tends to a positive constant as $u\to\infty$. We find $G(\sqrt{-\gamma})>0$, i.e. $G$ has no zeros.
Finally, if $m=-1$ imposing the conditions from Lemma 4.1 again leads to $x_{1}=-x_{2}$ with similar arguments as above, indeed, the only part that requires modification is the $\lambda>0,\gamma<0$ case. ∎
Apart from establishing the even parity of $B$ the argument above gives the following.
Corollary 4.6.
Let $x_{1},x_{2}$ with $x_{1}<x_{2}$ be the adjacent zeros of $B$ as in Lemma 4.1. Then $x_{1}=-x_{2}$ and in particular $B(0)>0$.
Proposition 4.7.
The function $B$ given by (1.3) and $b=0$ has two simple zeros and is positive between these zeros iff
$\lambda=0$ $\lambda>0$ $\lambda<0$
$m>0$ $c>0$ $c>\frac{\lambda}{m+1}$ $c>\frac{|\lambda|}{m+1}c_{0}$
$m\in(-1,0)$ $c>\frac{\lambda}{m+1}$
$m=-1$ $c>0$
$m<-1$ $c\in\Big{(}\frac{\lambda}{m+1},0\Big{)}$
where [^1] The asymptotic formula (4.6) implies that the minimum exists.
$c_{0}=\min_{x>x_{0}}\frac{(x^{2}+1)^{\frac{m}{2}+1}}{|F(x)|}$
and $x_{0}$ is the unique positive zero of $F$.
Proof.
We shall prove Proposition 4.7 by analysing all cases separately. It is convenient to introduce $\hat{B}(x):=(1+x^{2})^{m/2}B(x)$ which has the same zeros and sign as $B$.
Case 1. $m>0,\lambda\geq 0$
From the formula (1.3) we have $B(0)=c-\lambda/(m+1)$, so by Corollary 4.6
$c>\frac{\lambda}{m+1}.$
In particular, $c>0$. The zeros of $B$ coincide with those of
$\hat{B}(x)=cF(x)-\frac{\lambda}{(m+1)}(x^{2}+1)^{\frac{m}{2}+1}.$
Note that $\hat{B}(0)>0$, and $\hat{B}(x)\leq cF(x)$ for all $x>0$. Therefore, by Lemma 4.4, it follows that $B$ has precisely two roots, and $B$ is positive between these roots. A representative graph of $B$ is given in the figure below.
$m=3,c=1,\lambda=1$.
Case 2. $m>0,\lambda<0$
In this case
$B(0)=c+\frac{|\lambda|}{m+1}\;,$
and we require $B(0)>0$. The zeros of $B$ coincide with those of
$\hat{B}(x)=cF(x)+\frac{|\lambda|}{(m+1)}(x^{2}+1)^{\frac{m}{2}+1}.$
If $c\leq 0$ then $\hat{B}(x)$ for $x>0$ is a sum of increasing functions (Lemma 4.4) and since $\hat{B}(0)>0$ it follows that $B$ has no zeros. Now assume that $c>0$ which guarantees $B(0)>0$. Let $x_{0}$ be the unique positive zero of $F$, so that $F(x)<0$ for all $x>x_{0}$. To ensure that $B(x)<0$ for some $x>x_{0}$, and hence that $B$ has a zero for some $x>x_{0}$, it is necessary and sufficient that
$c>\frac{|\lambda|}{m+1}\min_{x>x_{0}}C(x),\quad\mbox{where}\quad C(x)=\frac{(x^{2}+1)^{\frac{m}{2}+1}}{|F(x)|}.$ (4.16)
The function $C(x)$ clearly has a minimum, as it is continuous for $x\in(x_{0},\infty)$, and (using the asymptotic form (4.6)) tends to $\infty$ as $x$ tends to $x_{0}$ or $\infty$. A representative graph of $B$ is given in the figure below.
$m=3,c=3,\lambda=-1$.
Finally we shall demonstrate that the two roots of $B$ are simple. Set $P=Bh$, where $h:(0,\infty)\rightarrow\mathbb{R}^{+}$. Then $B$ has a positive double zero at $x_{2}>0$ iff $P$ has a positive double zero at $x_{2}$. Now take $h=(x^{2}+1)^{m/2}/x$. Then $P^{\prime}(x)$ is given by the identity (4.7), and we find that it vanishes only at $x_{2}=(B(0)/|\lambda|)^{1/2}$. But $P$ tends to $\infty$ as $x$ tends to 00 and $\infty$, and we have picked a $c$ such that $P$ is negative somewhere. Therefore its minimum must be negative. So it does not have a double zero, and neither does $B$.
Case 3. $m<0,\lambda>0$
Reversing the arguments in the proof of Lemma 4.4 shows that $F(x)$ is monotonically increasing faster that a linear function, and in particular that it tends to $\infty$ from $F(0)=1$. There are now two cases to consider:
  • Case 3a. $m<-1$. The zeros of $B$ coincide with those of
    $\hat{B}(x)=cF(x)+\frac{\lambda}{|m+1|}(x^{2}+1)^{\frac{m}{2}+1},$ (4.17)
    and we have $\hat{B}(0)=B(0)>0$. If $c\geq 0$ then $cF$ is increasing, and since $\hat{B}(x)>cF(x)$ this implies $\hat{B}$ and hence $B$ has no zeros. Thus we assume that $c<0$. Therefore $cF(x)$ is decreasing faster than a linear function, and $cF(0)=c<0$. We have that $\hat{B}(0)>0$ iff
    $c\in\Big{(}\frac{\lambda}{m+1},0\Big{)}.$
    If $m<-1$ then $(x^{2}+1)^{\frac{m}{2}+1}<x$ for sufficiently large positive $x$. Therefore the first term in (4.17) is decreasing faster than a linear function, while the second term is either not increasing or increasing slower than $x^{1-\epsilon}$ for some $\epsilon>0$. Thus $\hat{B}\rightarrow-\infty$ as $x\rightarrow\infty$, and in particular it has two zeros. These are also the zeros of $B$. A representative graph of $B$ is given below.
    $m=-3,c=-1/5,\lambda=1$.
  • Case 3b. $m\in(-1,0)$. We have $B(0)>0$ iff
    $c>\frac{\lambda}{m+1}>0\;.$ (4.18)
    The zeros of $B$ coincide with those of
    $\hat{B}(x)=cF(x)-\frac{\lambda}{m+1}(1+x^{2})^{k},\quad\mbox{where}\quad k=\frac{m}{2}+1\in\Big{(}\frac{1}{2},1\Big{)}.$
    Using the asymptotic form (4.6) we see that the first term in $\hat{B}$ grows linearly for large $x$, whereas the second term in $\hat{B}$ decreases as $x^{m+2}$, and therefore $\hat{B}\to-\infty$ as $x\to\infty$. Therefore, $\hat{B}$ and hence $B$ has a zero. A representative graph of $B$ is given below.
    $m=-1/2,c=3,\lambda=1$.
Case 4. $m=-1,\lambda>0$
Consider the explicit formula for $B$ in (1.3) corresponding to $m=-1$, and take $c>0$. Therefore $F(0)=c>0$, and the existence two zeros follows as $x\;\operatorname{arcsinh}(x)$ grows faster that any constant multiple of $\sqrt{x^{2}+1}$. A representative graph is below.
$m=-1,c=1,\lambda=1$.
We are now ready to prove our main theorem.
Proof of Theorem 1.1. The Theorem now follows from the statements of Theorem 3.3, Proposition 4.7, and Lemmas 4.1 and 4.5 as follows: Theorem 3.3 gives the local normal form (1.2) and (1.3) under an assumption that $g$ has a Killing vector. Lemma 4.1 gives necessary and sufficient conditions on $B$ for the metric to extend to $S^{2}$, and Proposition 4.7 then shows that $B$ given by (1.3) with $b=0$ satisfies the conditions this Lemma. Finally all axi-symmetric solutions on $S^{2}$ arise from (1.2) and (1.3) with $b=0$ as a consequence of Lemma 4.5 which showed that $B$ must be even.
$\Box$
5. Projective metrizability and quasi–Einstein with $m=-1$
In this section we shall rule out the existence of non–trivial solutions to the quasi–Einstein equation (1.1) with $m=-1$ and $\lambda=0$ on any compact orientable surface. Proof of Theorem 1.2. Let $(g,X)$ satisfy the quasi–Einstein equation (1.1) on a surface $M$. Define an affine connection $D$ (possibly with torsion) by
$D=\nabla-p\;X^{\flat}\otimes\mbox{Id}-q\;\mbox{Id}\otimes X^{\flat}$ (5.1)
where $\nabla$ is the Levi–Civita connection of the metric $g$ and $(p,q)$ are constants. The torsion of this connection is $(p-q)\mbox{Id}\wedge X^{\flat}$, and the skew part of the Ricci tensor is $-\tfrac{1}{2}(2p+q)\text{d}X^{\flat}$. We verify by explicit calculation that the connection $D$ has vanishing Ricci tensor iff $m=-1,p=-\frac{1}{2},q=1$. We now refer to a result of Milnor [18] which states that if a closed surface admits a flat affine connection, then $M$ is diffeomorphic to a Klein bottle, or a torus. The Klein bottle is not orientable, which leaves $M=T^{2}$. It now follows from Proposition 2.1 that $X=0$ and the metric $g$ is flat.
$\Box$
The importance of this result stems from projective differential geometry [11], where it is an open problem to determine all affine connections with skew-symmetric Ricci tensor which share the unparametrised geodesics with a Levi–Civita connection of some (pseudo) Riemannian metric.
Definition 5.1.
A projective structure []delimited-[][$\nabla$][ ∇ ] is an equivalence class of torsion-free affine connections which share the same unparametrised geodesics. Two affine connections $\nabla$ and $\hat{\nabla}$ belong to the same projective structure iff there exists $\Upsilon\in\Lambda^{1}(M)$ such that
^=+$\otimes\Upsilon+\Upsilon\otimes\mbox{Id}.$ Id .
A projective structure is called:
  • metrisable if it contains a Levi–Civita connection of some metric in its projective class.
  • skew if it contains a representative connection with totally skew-symmetric Ricci tensor.
The Schouten tensor of a connection $\nabla$ is defined in terms of the Ricci tensor by
$\mathrm{P}_{ab}=\frac{1}{n-1}R_{(ab)}+\frac{1}{n+1}R_{[ab]},$
and is related to the Schouten tensor of $\hat{\nabla}$ by
$\hat{\mathrm{P}}_{ab}={\mathrm{P}}_{ab}-\nabla_{b}\Upsilon_{a}+\Upsilon_{a}\Upsilon_{b}.$ (5.2)
In case of two dimensional projective structures, the metrisability is obstructed by three projective differential invariants [2], which are differential expressions of order at least five in components of a chosen connection in the projective class. The skew projective structures have been partially characterised by projective differential invariants [21, 12, 19]. We aim to find metrisable connections projectively equivalent to skew connections. Assume that $\nabla$ is the Levi–Civita connection of some metric $g$, and seek a one–form $\Upsilon$ such that the Ricci tensor of $\hat{\nabla}$ is skew. Therefore $\hat{\mathrm{P}}_{(ab)}=0$, and the formula (5.2) with $\Upsilon=\frac{1}{1-n}X^{\flat}$ yields
$R_{ab}=-\frac{1}{n-1}X_{a}X_{b}-\nabla_{(a}X_{b)}$ (5.3)
which is the quasi–Einstein equation (1.1) with $m=1-n$ and $\lambda=0$. It is known [6] that compact surfaces admitting connections with skew–symmetric Ricci are diffeomorphic to a torus or a Klein bottle. Theorem 1.2 directly shows that the torus is necessarily flat if the connection is additionally metrisable.
Appendix
In [10] it was shown that the quasi Einstein-equations with $m=2$ on a closed $n$–dimensional manifold $M$, and with a non–gradient $X$ imply the existence of a Killing vector with a dual one–form $K^{\flat}=\text{d}\Gamma+\Gamma X^{\flat}$ for some smooth and positive function $\Gamma$. If additionally $n=2$ or $\lambda\leq 0$ then this Killing vector necessarily commutes with $X$. This result was sufficient to prove the rigidity of the extreme Kerr horizon with or without the cosmological constant. In this appendix we shall use the Fredholm alternative to show that $K$ commutes with $X$ in any dimension for any sign of the cosmological constant.
Theorem A.0.
Let $(M,g,X)$ be a compact quasi-Einstein manifold with $m=2$, and $X^{\flat}$ non-gradient. Suppose that $(M,g)$ admits a Killing vector field of the form $K^{\flat}=\mathrm{d}\Gamma+\Gamma X^{\flat}$ for some smooth and positive function $\Gamma$. Then $[K,X]=0$.
Proof.
Recall that (see Remark 2.6 in [10]) $[K,X]=0$ if and only if $\mathcal{L}_{K}\Gamma=0$. Consider the linear second order elliptic operator
$L\psi:=-\Delta\psi+\nabla_{a}((\Gamma^{-1}\nabla^{a}\Gamma)\psi)+\Gamma^{-2}|K|^{2}\psi$ (A.1)
with $\psi\in C^{\infty}(M)$. We will show that the kernel of $L$ is trivial and $L(\mathcal{L}_{K}\Gamma)=0$, which implies the result.
Lie deriving the equation $0=\nabla_{a}K^{a}=\nabla_{a}(\Gamma X^{a})+\Delta\Gamma$ along $K$ using the fact that $K$ is Killing gives
$0=\Delta(\mathcal{L}_{K}\Gamma)+\nabla_{a}(X^{a}\mathcal{L}_{K}\Gamma)+\nabla_{a}(\Gamma\mathcal{L}_{K}X^{a}).$ (A.2)
Lie deriving the trace of the quasi-Einstein equation (1.1) along $K$ shows that $\nabla_{a}(\mathcal{L}_{K}X^{a})=X_{a}\mathcal{L}_{K}X^{a}$. Combining with (A.2) we obtain
$0=\Delta(\mathcal{L}_{K}\Gamma)+X^{a}\nabla_{a}(\mathcal{L}_{K}\Gamma)+(\mathcal{L}_{K}\Gamma)\nabla_{a}X^{a}+K_{a}\mathcal{L}_{K}X^{a}.$ (A.3)
We can express $\mathcal{L}_{K}X^{a}$ in terms of $\mathcal{L}_{K}\Gamma$ using $X^{a}=\Gamma^{-1}(K^{a}-\nabla^{a}\Gamma)$. Explicitly,
$\mathcal{L}_{K}X^{a}=-\Gamma^{-1}(X^{a}\mathcal{L}_{K}\Gamma+\nabla^{a}(\mathcal{L}_{K}\Gamma)).$ (A.4)
Inserting (A.4) into (A.3) and writing $X^{a}$ in terms of $K^{a},\Gamma$ gives the equation $L(\mathcal{L}_{K}\Gamma)=0$ where $L$ is the operator defined above.
It remains to prove that ker $L=\{0\}$. By the Fredholm alternative the dimension of ker $L$ is finite and equal to the dimension of ker $L^{*}$, where $L^{*}$ is the formal adjoint of $L$ in the $L^{2}$ inner product. In the case at hand,
$L^{*}\psi=-\Delta\psi-(\Gamma^{-1}\nabla^{a}\Gamma)\nabla_{a}\psi+|K|^{2}\Gamma^{-2}\psi.$ (A.5)
This operator is of the form $L^{*}\psi=-\Delta\psi+B^{a}\nabla_{a}\psi+C\psi$ with $C\geq 0$. Note that $C=|K|^{2}\Gamma^{-2}$ is not identically zero because $X^{\flat}$ is non-gradient. By the strong maximum principle (Theorem 2.9 in [13]) it follows that the kernel of $L^{*}$ is trivial, which concludes the proof. ∎